A Region of land 50meters by 40meters is to cross diagonally by a road 15meters wide. What is the roads area?


1
Expert's answer
2022-03-14T10:30:11-0400



we start by calculating the value of x

"15^2=x^2+x^2"

"225=2x^2"

"225\/2=x^2"

"112.5=x^2"

"\sqrt{112.5}= \sqrt{x^2}"

"10.61=x"

length x= 10.61m

next we find the diagonal D

length = "50-x=50-10.61=39.39"

width ="40-x=40-10.61=29.39"

D="\sqrt{39.39^2+29.39^2}"

D="49.15"

area of the road ="15*D"

="15*49.15"

="737.25m^2"